deanna is five more than six times than ted. Twine the sum of their ages is 94. How old is ted?
Answer by elima(1433)

deanna is five more than six times than ted. Twine the sum of their ages is 94. How old is ted?
x= Deana's age
y= Ted's age
=============
x=6y+5
2x+2y=94
===============
substitute x=6y+5 for x in second equation
2(6y+5)=94
12y+10=94
12y=84
y=6
Now plug that in the first equation for y;
x=6(6)+5
x=36+5
x=41
:)